OpenStudy (anonymous):

Rewrite in simplest rational exponent form √x • 4√x. Please show each step of your process

zepdrix (zepdrix):

Hey King! Is this what your problem looks like?\[\large\rm \sqrt{x}\cdot\sqrt[4]{x}\]

OpenStudy (anonymous):

yes :)

zepdrix (zepdrix):

When no number is written over the radical, it's implied that it's a 2.\[\large\rm \sqrt[2]{x}\cdot\sqrt[4]{x}\]Square root = second root. Ok so in general,\[\large\rm \sqrt[n]{x}=x^{1/n}\]We can rewrite our root as a rational exponent like this. We get a fraction as an exponent. The `denominator` is the degree of the root.

zepdrix (zepdrix):

\[\large\rm \sqrt[2]{x}\cdot\sqrt[4]{x}\quad=\quad x^{1/2}\cdot \text{__}\] So that's a way I could rewrite my first radical. How bout the other one? :) Any ideas?

OpenStudy (anonymous):

would you reduce the second one to 2?

zepdrix (zepdrix):

No.\[\LARGE\rm \sqrt[\color{orangered}{n}]{x}=x^{1/\color{orangered}{n}}\qquad\to\qquad \sqrt[\color{orangered}{4}]{x}=x^{1/\color{orangered}{4}}\]

zepdrix (zepdrix):

\[\large\rm \sqrt[2]{x}\cdot\sqrt[4]{x}\quad=\quad x^{1/2}\cdot x^{1/4}\]

zepdrix (zepdrix):

From here you need to apply one of your exponent rules:\[\large\rm \color{royalblue}{x^a\cdot x^b=x^{a+b}}\]

OpenStudy (anonymous):

oh yeah sorry i was confused but I see how it's 1/4

OpenStudy (anonymous):

@zepdrix is that it after that step ?

zepdrix (zepdrix):

You need to apply the blue formula. So you would have to add 1/2 and 1/4.

zepdrix (zepdrix):

You'll need a common denominator to add those two fractions together.

OpenStudy (anonymous):

the common denominator would be 2 right ?

OpenStudy (anonymous):

@zepdrix

zepdrix (zepdrix):

\[\large\rm \frac{1}{2}+\frac{1}{4}\]The common denominator is 4. How do you turn the 2 into a 4?

OpenStudy (anonymous):

multiplying it by 2

zepdrix (zepdrix):

\[\large\rm \color{royalblue}{\frac{2}{2}}\cdot\frac{1}{2}+\frac{1}{4}\quad=\quad \frac{2}{4}+\frac{1}{4}\]Good good good.

OpenStudy (anonymous):

wait when you said I had to apply one of the exponent rules what was it called?

zepdrix (zepdrix):

It was the one that I colored in blue. I think it's called the Exponent Addition Rule

zepdrix (zepdrix):

something like that

OpenStudy (anonymous):

the final answer simplified and all would be x^3/4 or am I wrong ? @zepdrix

zepdrix (zepdrix):

yay good job \c:/

OpenStudy (anonymous):

thank you for the help. it was much needed